# arXiv:2602.24042v1[cs.DS] 27 Feb 2026

### v1

@@ A.2 Additional Related Work
Some previous work showed that given $O(\varepsilon)$ resource augmentation, there are $PTAS$ for approximating the optimal adaptive and non-adaptive policies for $\operatorname{Non-Risky-SK}$ and $\operatorname{Risky-SK}$.
[BGK11] give a $PTAS$ for computing adaptive and non-adaptive policies for $\operatorname{Non-Risky-SK}$ and $\operatorname{Risky-SK}$. [BGK11] introduce and use the notion of block-adaptive policies (policies deciding on blocks of items to insert and not just individual items). They show that (1) there exists a $1+\varepsilon$ adaptive policy that inserts at most $O(poly(\frac{1}{\varepsilon}))$ blocks, and (2) enumerating all of the possible block-adaptive policies is possible in $\varepsilon$-dependent polynomial time. However, their policy construction requires querying the state of the knapsack for each item in each inserted block, leading to a linear number of adaptive queries for the state of the knapsack. Also, their policies require a large ($O(\frac{1}{\varepsilon^{15}})$) number of blocks, and more importantly, $O(\varepsilon)$ resource augmentation. The same principle applies to the PTAS of [LY13] for $\operatorname{Non-Risky-SK}$ and of [FLX18] to $\operatorname{Risky-SK}$, which both follow the same idea of searching a close to optimal block-adaptive policies [BGK11] (while their exact methods for searching the block-adaptive policy space differ). All of these PTAS results require both $O(\varepsilon)$ resource augmentation, and knowing (thus querying) the state of the knapsack inside each block to decide on the next block to insert. These results are not fully polynomial and the dependence on $\varepsilon$ implies a large run time run time in practice. In this work, we focus on policies that (1) have no resource augmentation, (2) use only at most $\tilde{O}\left(\frac{1}{\varepsilon}\right)$ adaptive queries, and (3) query only the current state of the knapsack rather than the entire history of realizations.
@@
[Gup+11] show an $8$ approximate adaptive policy for the variant of the problem where cancellations (throwing an item away) and correlations (between item sizes and values) are allowed. This was later improved by [Ma14] who provides $2-\varepsilon$ approximate (fully adaptive) policy by introducing additional assumptions (they assume that if the knapsack capacity is $T$, the processing times of each jobs are integers. Also, the algorithm run time is polynomial in $T$).
@@
Another studied variant is the $p$-Chance Constrained Stochastic Knapsack, which is also known as the bounded overflow probability model. The goal in this variant is to maximize the value under the constraint that the overflow probability is at most $p$. In this bounded overflow probability model, [KRT97] give a $log(1/\gamma)$- factor approximation algorithms when items sizes have Bernoulli size distribution. [GI99] give PTAS and QPTAS results for restricted classes of distributions, namely Bernoulli, Exponential and Poisson distribution. [De18] provide approximation schemes (that do not require $1+O(\varepsilon)$ resource augmentation) for the cases where the item sizes are Bernoulli, k-supported or hyper-contractive random variables. These versions are just like $\operatorname{Risky-SK}$, except overflow probability $p$ is allowed (the algorithm gets the value as long as the probability for overflow is at most $p$). We note that it is not true that any policy for $\operatorname{Risky-SK}$ if also a policy for $\varepsilon$-chance constrained problem for any small $\varepsilon$ — as Example 1 demonstrates, the optimal adaptive policy may have probability as high as $\frac{1}{e}$ of overflow. One can show that the optimal non-adaptive policy for the same instance also has a constant positive probability of overflow.
@@ B.1 Proof of 2
(b) Large $\beta$, small $p$. If $\beta\geq\beta_{v}$ and $p<p_{\min}(\beta)$, then $B\geq\frac{1}{4}\geq v$ as we now show:
@@
(c) Large $\beta$, large $p$. If $\beta\geq\beta_{v}$ and $p\geq p_{v}$, then
@@
    For any fixed $\beta$, let $G_{\beta}(p):=(1-p)\beta(1-\beta+p)=-\beta p^{2}+\beta^{2}p+\beta(1-\beta)$. $G_{\beta}(p)$ is a concave quadratic function in $p$. Hence its minimum on any interval is attained at an endpoint of that interval. Plugging in the left endpoint $p_{\min}=\frac{4\beta-2}{\,4\beta-1\,}=1-\frac{1}{4\beta-1}$ yields
@@ C.2 Proof of 3
So the algorithm that partitions $I$ into $I_{1}$, $I_{2}$ and either runs $A_{1}$ on $I_{1}$ or $A_{2}$ on $I_{2}$ according to the option with the maximum expected value gets an expected value of at least
@@ D.1 Proof of 7
We show the proof for $H_{2}$, the proof for $H_{2}^{T}$ is identical: any tree constraint makes sure one can only choose root-to-leaf path, but so are the size distribution choices of $H_{2}^{\prime}$. Let any instance $I\in\mathcal{H}_{2}$. Then $\operatorname{\sigma_{AD}}$ gets its value from taking item 0 and then some other item $i$. First, we show the first property. That is, we show a transformation from instance $I$ to instance $I^{\prime}$ where we replace each item with a set of deterministic items (we say an item is deterministic if it has a deterministic size distribution).
@@
where the second inequality is due to the definition of $v_{i,c_{i}}$ and the last inequality is since in $I^{\prime}$ after taking item 0 and seeing $S_{0}=c_{i}$ the adaptive optimum may choose to take item $i,c_{i}$} which fits the knapsack w.p. $1$.
The intuition is that for every realization of $S_{0}$ to size $c$ in instance $I$, there is a choice of next item to insert in $I^{\prime}$ that yields at least the same value and always fits the knapsack.
@@
For any $j\in[n],c\in C$, Let $\operatorname{ALG}_{j}$, $\operatorname{ALG}^{\prime}_{j,c}$ denote the value the best non-adaptive policy gets from choosing item 0 and then item $j$ in instance $I$ and the value it gets from choosing item 0 and then item $j,c$ in instance $I^{\prime}$, respectively.
We now show that for any new such option for $\operatorname{\sigma_{ALG}}$ in $I^{\prime}$ yields less value than the original option for $\operatorname{\sigma_{ALG}}$ in $I$:
@@
For any $i\neq 0$ the option of choosing item $i$ and then item 0 yields the same value for as choosing item 0 and then item $i$ (due to order invariance). Eq. 30 means that any such option yields less value in $I^{\prime}$ than in $I$. Let us rename the items of $I^{\prime}$ by their value such that $v_{1}\geq...\geq v_{n^{\prime}}$. We can assume w.l.o.g. that $S_{i}>S_{j}$ for any $i<j$ as otherwise no policy would ever choose item $j$ over item $i$.
@@
We conclude: for any non-adaptive policy for $I^{\prime}$ there’s a non-adaptive policy for $I$ that yields at least the same value: $\operatorname{ALG}(I^{\prime})\leq\operatorname{ALG}(I)$.
@@ D.2 Proof of 8
For the base: consider an instance of size 2. So the policy only has two options: either take item 0 and then item $1$, or take item $1$ (order-invariance implies that taking item $1$ and then item 0 yield the same value). If taking item $1$ yields more value for $\operatorname{\sigma_{ALG}}$: $v_{1}>(1-p_{0})p_{1}(1+v_{1})$, then we could decrease $p_{0}$ and increase $p_{1}$ until the two options are equal (from intermediate value theorem). $\operatorname{ADAPT}$ only increases while $\operatorname{ALG}$ stays the same. If $v_{1}<(1-p_{0})p_{1}(1+v_{1})$ then we increase $p_{0}$ until taking item $1$ and stopping yields the same value.
@@
Case (1): Assume taking item $1$ yields the (strictly) best expected value for $\operatorname{\sigma_{ALG}}$: $v_{1}>(1-p_{0})p_{1}(1+v_{1})$: just like in the base case, we could always decrease $p_{0}$ and increase $p_{1}$ (while decreasing the probability mass from the rest of the options of $S_{0}$) until the two terms are equal. As before intermediate value theorem implies there are such $p_{0}$$p_{1}$ values such that the two terms are equal. In the new resulting instance, $I^{\prime}_{1}$, $ADPAT$ increases (as we shift probability from lower valued items to a bigger valued item) and $\operatorname{ALG}$ does not change.
Next, we have a series of transformations from $I^{\prime}_{i}$ to $I^{\prime}_{i+1}$ for any $i\in[n-1]$: We increase the value $v_{i+1}$ until either (1a) it equals $v_{i}$ or (1b) choosing item 0 and then item $i+1$ yields the same expected value as choosing (only) item $1$. That is: $v_{1}=(1-p_{0})\sum_{j=1}^{i}p_{j}(1+v_{i+1})$.
@@
In case (1b), via an additional simple induction, we get that all of the options of $\operatorname{\sigma_{ALG}}$ (of choosing item $1$ only, or choosing item 0 and then item $j\in\left\{1,\ldots,i+1\right\}$) yield the same expected value. The process stops when either an instance of n1n-1 items with the same adaptivity gap is obtained (and can therefore use the induction hypothesis) or when we reach $I^{\prime}_{n}$ and get that all options yield the same expected value, as required.
Case (2): Taking item 0 and then item $1$ is the (strictly) best option. In this case, we can decrease $p_{1}$ and increase $p_{0}$ by the same amount until this is no longer the case. From intermediate value theorem there are such $p_{0},p_{1}$ values. The adaptivity gap may only increase due to the change: Let $\operatorname{ADAPT}^{\prime}$$\operatorname{ALG}^{\prime}$ be the expected value of $\operatorname{\sigma_{AD}}$ and $\operatorname{\sigma_{ALG}}$ in the new resulting instance. The change in $p_{0}$ translates into a multiplicative change in $\operatorname{ADAPT}$, $\operatorname{ALG}$ that affects both of them in the same way and so the ratio between the two is not affected by a change in $p_{0}$. The change in $p_{1}$. Let $\Delta=(1-p_{0})(p_{1}-p^{\prime}_{1})(v_{1}+1)$. Then:
@@
The adaptivity gap has only increased: $G(I^{\prime})\geq G(I)$. After this change, the option of taking only item $1$ and the option of taking item 0 and then item $1$ both yield the same expected value, which is the value of $\operatorname{ALG}$. We thus we may continue as in case (1).
Case (3): If the (strictly) best option of $\operatorname{\sigma_{ALG}}$ is to take item 0 and then some item $i>1$, then $v_{1}<(1-p_{0})(\sum_{j=1}^{i}p_{j})(1+v_{i})$. We could decrease $p_{i}$ and increase $p_{1}$ by the same amount until finally either (a) $p_{i}=0$ (and therefore no optimal policy, adaptive or not, will choose item $i$) or (b) via intermediate value theorem we get that the option of choosing item 0 and then item $1$ yields the same value as the option of choosing item 0 and then item $i$.
@@
Case (3b): In the second case, the two options yield the same expected value (while the adaptivity gap may only increase). We continue this way for all $i>1$ so there is no $i>1$ such that taking it after item 0 yields the strictly best option of $\operatorname{\sigma_{ALG}}$. In the resulting instance: either taking only item $1$ or taking item 0 and then taking item $1$ is the best option. If taking only item $1$ is strictly the better option, and we can continue according to case (1) or case (2).
@@ D.3 Proof of 12
We start with the first property: Let $I$ be an instance in $\mathcal{H}^{\prime}_{2}$. First, we can obtain an instance $I^{\prime}$ by replacing each item $i\in[n]$ with a new item such that the value of the new item $v^{\prime}_{i}=a\cdot v_{i}$ and its size $S^{\prime}_{i}$ is $S_{i}$ w.p. $\frac{1}{a}$ and $2$ with probability $1-\frac{1}{a}$. In expectation, every policy (adaptive or not) gets the same value for item $i$ and for item $i^{\prime}$. The difference is that now if a policy picks item $i^{\prime}$ first, it would gain almost nothing ($\frac{1}{a}\ll 1$) by choosing any other item later on. $\operatorname{ADAPT}$ may only increase due to this change but $\operatorname{ALG}$ can only decrease (all options yield the same expected value except for ones starting with item $i\neq 0$ and then picking some other item). So indeed $G(I^{\prime})\geq G(I)-O(\frac{1}{a})$.
@@ D.4 Proof of 13
Throughout the proof when we refer to $v_{i}$ as its effective value rather than value for any $i\in[n]$ (otherwise we could use $w_{i}=\frac{1}{a}v_{i}$ instead). The base case is for $n=1$. In this case there are two items: item 0 and $1$. Consider the case where taking item $1$ yields higher value than taking item 0 and then $1$. If $v_{1}>(1-p_{0})(1+p_{1}v_{1})$ we could always decrease $p_{0}$ until either the two sides are the same or until $p_{0}=0$ (via intermediate value theorem). In the second option $v_{1}>1+p_{1}v_{1}$. We could increase $p_{1}$ such that $v_{1}=1+p_{1}v_{1}$. There is always such a $p_{1}$ (from intermediate value theorem) since $f(p_{1})=v_{1}-(1+p_{1}v_{1})$ is positive for the initial $p_{1}$ but is negative for $p_{1}=1$. Hence, in the new instance we obtain, the two options of $\operatorname{\sigma_{ALG}}$ yield the same expected value. Similarly, if taking item 0 and then $1$ yields higher value we could increase $p_{0}$ until the two options yield the same expected value.
@@
Case (1): Assume taking item $1$ yields the (strictly) best expected value for $\operatorname{\sigma_{ALG}}$: $v_{1}>(1-p_{0})(1+p_{1}v_{1})$: just like in the base case, we could always decrease $p_{0}$ and increase $p_{1}$ (while decreasing the probability mass from the rest of the options of $S_{0}$) until the two terms are equal. As before intermediate value theorem implies there are such $p_{0}$$p_{1}$ values such that the two terms are equal. In the new resulting instance, $I^{\prime}_{1}$, $ADPAT$ increases (as we shift probability from lower valued items to a bigger valued item) and $\operatorname{ALG}$ may only decrease.
Next, we have a series of transformations from $I^{\prime}_{i}$ to $I^{\prime}_{i+1}$ for any $i\in[n-1]$: We increase the value $v_{i+1}$ until either (1a) it equals $v_{i}$ or (1b) choosing item 0 and then item $i+1$ yields the same expected value as choosing item $1$. That is: $v_{1}=(1-p_{0})(1+\sum_{j=1}^{i}p_{j}v_{i+1})$.
@@
In case (1b), via an additional simple induction, we get that all of the options of $\operatorname{\sigma_{ALG}}$ (of choosing item $1$ only, or choosing item 0 and then item $j\in\left\{1,\ldots,i+1\right\}$) yield the same expected value. The process stops when either an instance of n1n-1 items with the same adaptivity gap is obtained (and can therefore use the induction hypothesis) or when we reach $I^{\prime}_{n}$ and get that all options yield the same expected value, as required.
Case (2): Taking item 0 and then item $1$ is the (strictly) best option. In this case, we can decrease $p_{1}$ and increase $p_{0}$ by the same amount until this is no longer the case. From intermediate value theorem there is such a $p_{1}$. The adaptivity gap may only increase due to the change: Let $\operatorname{ADAPT}^{\prime}$$\operatorname{ALG}^{\prime}$ be the expected value of $\operatorname{\sigma_{AD}}$ and $\operatorname{\sigma_{ALG}}$ in the new resulting instance. Let $\Delta=(1-p_{0})(p_{1}-p^{\prime}_{1})v_{1}$ be the difference between $\operatorname{ADAPT}$ and $\operatorname{ADAPT}^{\prime}$, and similarly between $\operatorname{ALG}$ and $\operatorname{ALG}^{\prime}$. Then $\frac{\operatorname{ADAPT}^{\prime}}{\operatorname{ALG}^{\prime}}=\frac{\operatorname{ADAPT}-\Delta}{\operatorname{ALG}-\Delta}\geq\frac{\operatorname{ADAPT}}{\operatorname{ALG}}$; the adaptivity gap has only increased. After the change the option of taking only item $1$ and the option of taking item 0 and then item $1$ both yield the same expected value, which is the value of $\operatorname{ALG}$. We thus may continue as in case (1).
Case (3): If the (strictly) best option of $\operatorname{\sigma_{ALG}}$ is to take item 0 and then some item $i>1$, then $v_{1}<(1-p_{0})(1+\sum_{j=1}^{i}p_{j}v_{i})$. We could decrease $p_{i}$ and increase $p_{1}$ by the same amount until finally either (a) $p_{i}=0$ (and therefore no optimal policy, adaptive or not, will choose item $i$) or (b) via intermediate value theorem we get that the option of choosing item 0 and then item $1$ yields the same value as the option of choosing item 0 and then item $i$.
@@
Case (3b): In the second case, the two options yield the same expected value (while the adaptivity gap may only increase). We continue this way for all $i>1$ so there is no $i>1$ such that taking it after item 0 yields the strictly best option of $\operatorname{\sigma_{ALG}}$. In the resulting instance: either taking only item $1$ or taking item 0 and then taking item $1$ is the best option. If taking only item $1$ is strictly the better option, and we can continue according to case (1) or case (2).

### v2

@@ A.2 Additional Related Work
Some previous work showed that given $O(\varepsilon)$ resource augmentation, there are PTASPTAS for approximating the optimal adaptive and non-adaptive policies for $\operatorname{Non-Risky-SK}$ and $\operatorname{Risky-SK}$.
[BGK11] give a PTASPTAS for computing adaptive and non-adaptive policies for $\operatorname{Non-Risky-SK}$ and $\operatorname{Risky-SK}$. [BGK11] introduce and use the notion of block-adaptive policies (policies deciding on blocks of items to insert and not just individual items). They show that (1) there exists a $1+\varepsilon$ adaptive policy that inserts at most $O(poly(\frac{1}{\varepsilon}))$ blocks, and (2) enumerating all of the possible block-adaptive policies is possible in $\varepsilon$-dependent polynomial time. However, their policy construction requires querying the state of the knapsack for each item in each inserted block, leading to a linear number of adaptive queries for the state of the knapsack. Also, their policies require a large ($O(\frac{1}{\varepsilon^{15}})$) number of blocks, and more importantly, $O(\varepsilon)$ resource augmentation. The same principle applies to the PTAS of [LY13] for $\operatorname{Non-Risky-SK}$ and of [FLX18] to $\operatorname{Risky-SK}$, which both follow the same idea of searching a close to optimal block-adaptive policies [BGK11] (while their exact methods for searching the block-adaptive policy space differ). All of these PTAS results require both $O(\varepsilon)$ resource augmentation, and knowing (thus querying) the state of the knapsack inside each block to decide on the next block to insert. These results are not fully polynomial and the dependence on $\varepsilon$ implies a large run time run time in practice. In this work, we focus on policies that (1) have no resource augmentation, (2) use only at most $\tilde{O}\left(\frac{1}{\varepsilon}\right)$ adaptive queries, and (3) query only the current state of the knapsack rather than the entire history of realizations.
@@
[Gup+11] show an 88 approximate adaptive policy for the variant of the problem where cancellations (throwing an item away) and correlations (between item sizes and values) are allowed. This was later improved by [Ma14] who provides $2-\varepsilon$ approximate (fully adaptive) policy by introducing additional assumptions (they assume that if the knapsack capacity is $T$, the processing times of each jobs are integers. Also, the algorithm run time is polynomial in $T$).
@@
Another studied variant is the $p$-Chance Constrained Stochastic Knapsack, which is also known as the bounded overflow probability model. The goal in this variant is to maximize the value under the constraint that the overflow probability is at most pp. In this bounded overflow probability model, [KRT97] give a $log(1/\gamma)$- factor approximation algorithms when items sizes have Bernoulli size distribution. [GI99] give PTAS and QPTAS results for restricted classes of distributions, namely Bernoulli, Exponential and Poisson distribution. [De18] provide approximation schemes (that do not require $1+O(\varepsilon)$ resource augmentation) for the cases where the item sizes are Bernoulli, k-supported or hyper-contractive random variables. These versions are just like $\operatorname{Risky-SK}$, except overflow probability $p$ is allowed (the algorithm gets the value as long as the probability for overflow is at most $p$). We note that it is not true that any policy for $\operatorname{Risky-SK}$ if also a policy for $\varepsilon$-chance constrained problem for any small $\varepsilon$ — as Example 1 demonstrates, the optimal adaptive policy may have probability as high as $\frac{1}{e}$ of overflow. One can show that the optimal non-adaptive policy for the same instance also has a constant positive probability of overflow.
@@ B.1 Proof of 2
(b) Large $\beta$, small pp. If $\beta\geq\beta_{v}$ and $p<p_{\min}(\beta)$, then $B\geq\frac{1}{4}\geq v$ as we now show:
@@
(c) Large $\beta$, large pp. If $\beta\geq\beta_{v}$ and $p\geq p_{v}$, then
@@
    For any fixed $\beta$, let $G_{\beta}(p):=(1-p)\beta(1-\beta+p)=-\beta p^{2}+\beta^{2}p+\beta(1-\beta)$. $G_{\beta}(p)$ is a concave quadratic function in pp. Hence its minimum on any interval is attained at an endpoint of that interval. Plugging in the left endpoint $p_{\min}=\frac{4\beta-2}{\,4\beta-1\,}=1-\frac{1}{4\beta-1}$ yields
@@ C.2 Proof of 3
So the algorithm that partitions II into $I_{1}$, $I_{2}$ and either runs $A_{1}$ on $I_{1}$ or $A_{2}$ on $I_{2}$ according to the option with the maximum expected value gets an expected value of at least
@@ D.1 Proof of 7
We show the proof for $H_{2}$, the proof for $H_{2}^{T}$ is identical: any tree constraint makes sure one can only choose root-to-leaf path, but so are the size distribution choices of $H_{2}^{\prime}$. Let any instance $I\in\mathcal{H}_{2}$. Then $\operatorname{\sigma_{AD}}$ gets its value from taking item 0 and then some other item ii. First, we show the first property. That is, we show a transformation from instance II to instance $I^{\prime}$ where we replace each item with a set of deterministic items (we say an item is deterministic if it has a deterministic size distribution).
@@
where the second inequality is due to the definition of $v_{i,c_{i}}$ and the last inequality is since in $I^{\prime}$ after taking item 0 and seeing $S_{0}=c_{i}$ the adaptive optimum may choose to take item $i,c_{i}$} which fits the knapsack w.p. 11.
The intuition is that for every realization of $S_{0}$ to size $c$ in instance II, there is a choice of next item to insert in $I^{\prime}$ that yields at least the same value and always fits the knapsack.
@@
For any $j\in[n],c\in C$, Let $\operatorname{ALG}_{j}$, $\operatorname{ALG}^{\prime}_{j,c}$ denote the value the best non-adaptive policy gets from choosing item 0 and then item $j$ in instance II and the value it gets from choosing item 0 and then item $j,c$ in instance $I^{\prime}$, respectively.
We now show that for any new such option for $\operatorname{\sigma_{ALG}}$ in $I^{\prime}$ yields less value than the original option for $\operatorname{\sigma_{ALG}}$ in II:
@@
For any $i\neq 0$ the option of choosing item $i$ and then item 0 yields the same value for as choosing item 0 and then item $i$ (due to order invariance). Eq. 30 means that any such option yields less value in $I^{\prime}$ than in II. Let us rename the items of $I^{\prime}$ by their value such that $v_{1}\geq...\geq v_{n^{\prime}}$. We can assume w.l.o.g. that $S_{i}>S_{j}$ for any $i<j$ as otherwise no policy would ever choose item $j$ over item ii.
@@
We conclude: for any non-adaptive policy for $I^{\prime}$ there’s a non-adaptive policy for II that yields at least the same value: $\operatorname{ALG}(I^{\prime})\leq\operatorname{ALG}(I)$.
@@ D.2 Proof of 8
For the base: consider an instance of size 2. So the policy only has two options: either take item 0 and then item 11, or take item 11 (order-invariance implies that taking item 11 and then item 0 yield the same value). If taking item 11 yields more value for $\operatorname{\sigma_{ALG}}$: $v_{1}>(1-p_{0})p_{1}(1+v_{1})$, then we could decrease $p_{0}$ and increase $p_{1}$ until the two options are equal (from intermediate value theorem). $\operatorname{ADAPT}$ only increases while $\operatorname{ALG}$ stays the same. If $v_{1}<(1-p_{0})p_{1}(1+v_{1})$ then we increase $p_{0}$ until taking item 11 and stopping yields the same value.
@@
Case (1): Assume taking item 11 yields the (strictly) best expected value for $\operatorname{\sigma_{ALG}}$: $v_{1}>(1-p_{0})p_{1}(1+v_{1})$: just like in the base case, we could always decrease $p_{0}$ and increase $p_{1}$ (while decreasing the probability mass from the rest of the options of $S_{0}$) until the two terms are equal. As before intermediate value theorem implies there are such $p_{0}$$p_{1}$ values such that the two terms are equal. In the new resulting instance, $I^{\prime}_{1}$, ADPATADPAT increases (as we shift probability from lower valued items to a bigger valued item) and $\operatorname{ALG}$ does not change.
Next, we have a series of transformations from $I^{\prime}_{i}$ to $I^{\prime}_{i+1}$ for any $i\in[n-1]$: We increase the value $v_{i+1}$ until either (1a) it equals $v_{i}$ or (1b) choosing item 0 and then item $i+1$ yields the same expected value as choosing (only) item 11. That is: $v_{1}=(1-p_{0})\sum_{j=1}^{i}p_{j}(1+v_{i+1})$.
@@
In case (1b), via an additional simple induction, we get that all of the options of $\operatorname{\sigma_{ALG}}$ (of choosing item 11 only, or choosing item 0 and then item $j\in\left\{1,\ldots,i+1\right\}$) yield the same expected value. The process stops when either an instance of n1n-1 items with the same adaptivity gap is obtained (and can therefore use the induction hypothesis) or when we reach $I^{\prime}_{n}$ and get that all options yield the same expected value, as required.
Case (2): Taking item 0 and then item 11 is the (strictly) best option. In this case, we can decrease $p_{1}$ and increase $p_{0}$ by the same amount until this is no longer the case. From intermediate value theorem there are such $p_{0},p_{1}$ values. The adaptivity gap may only increase due to the change: Let $\operatorname{ADAPT}^{\prime}$$\operatorname{ALG}^{\prime}$ be the expected value of $\operatorname{\sigma_{AD}}$ and $\operatorname{\sigma_{ALG}}$ in the new resulting instance. The change in $p_{0}$ translates into a multiplicative change in $\operatorname{ADAPT}$, $\operatorname{ALG}$ that affects both of them in the same way and so the ratio between the two is not affected by a change in $p_{0}$. The change in $p_{1}$. Let $\Delta=(1-p_{0})(p_{1}-p^{\prime}_{1})(v_{1}+1)$. Then:
@@
The adaptivity gap has only increased: $G(I^{\prime})\geq G(I)$. After this change, the option of taking only item 11 and the option of taking item 0 and then item 11 both yield the same expected value, which is the value of $\operatorname{ALG}$. We thus we may continue as in case (1).
Case (3): If the (strictly) best option of $\operatorname{\sigma_{ALG}}$ is to take item 0 and then some item $i>1$, then $v_{1}<(1-p_{0})(\sum_{j=1}^{i}p_{j})(1+v_{i})$. We could decrease $p_{i}$ and increase $p_{1}$ by the same amount until finally either (a) $p_{i}=0$ (and therefore no optimal policy, adaptive or not, will choose item ii) or (b) via intermediate value theorem we get that the option of choosing item 0 and then item 11 yields the same value as the option of choosing item 0 and then item ii.
@@
Case (3b): In the second case, the two options yield the same expected value (while the adaptivity gap may only increase). We continue this way for all $i>1$ so there is no $i>1$ such that taking it after item 0 yields the strictly best option of $\operatorname{\sigma_{ALG}}$. In the resulting instance: either taking only item 11 or taking item 0 and then taking item 11 is the best option. If taking only item 11 is strictly the better option, and we can continue according to case (1) or case (2).
@@ D.3 Proof of 12
We start with the first property: Let II be an instance in $\mathcal{H}^{\prime}_{2}$. First, we can obtain an instance $I^{\prime}$ by replacing each item $i\in[n]$ with a new item such that the value of the new item $v^{\prime}_{i}=a\cdot v_{i}$ and its size $S^{\prime}_{i}$ is $S_{i}$ w.p. $\frac{1}{a}$ and 22 with probability $1-\frac{1}{a}$. In expectation, every policy (adaptive or not) gets the same value for item $i$ and for item $i^{\prime}$. The difference is that now if a policy picks item $i^{\prime}$ first, it would gain almost nothing ($\frac{1}{a}\ll 1$) by choosing any other item later on. $\operatorname{ADAPT}$ may only increase due to this change but $\operatorname{ALG}$ can only decrease (all options yield the same expected value except for ones starting with item $i\neq 0$ and then picking some other item). So indeed $G(I^{\prime})\geq G(I)-O(\frac{1}{a})$.
@@ D.4 Proof of 13
Throughout the proof when we refer to $v_{i}$ as its effective value rather than value for any $i\in[n]$ (otherwise we could use $w_{i}=\frac{1}{a}v_{i}$ instead). The base case is for $n=1$. In this case there are two items: item 0 and 11. Consider the case where taking item 11 yields higher value than taking item 0 and then 11. If $v_{1}>(1-p_{0})(1+p_{1}v_{1})$ we could always decrease $p_{0}$ until either the two sides are the same or until $p_{0}=0$ (via intermediate value theorem). In the second option $v_{1}>1+p_{1}v_{1}$. We could increase $p_{1}$ such that $v_{1}=1+p_{1}v_{1}$. There is always such a $p_{1}$ (from intermediate value theorem) since $f(p_{1})=v_{1}-(1+p_{1}v_{1})$ is positive for the initial $p_{1}$ but is negative for $p_{1}=1$. Hence, in the new instance we obtain, the two options of $\operatorname{\sigma_{ALG}}$ yield the same expected value. Similarly, if taking item 0 and then 11 yields higher value we could increase $p_{0}$ until the two options yield the same expected value.
@@
Case (1): Assume taking item 11 yields the (strictly) best expected value for $\operatorname{\sigma_{ALG}}$: $v_{1}>(1-p_{0})(1+p_{1}v_{1})$: just like in the base case, we could always decrease $p_{0}$ and increase $p_{1}$ (while decreasing the probability mass from the rest of the options of $S_{0}$) until the two terms are equal. As before intermediate value theorem implies there are such $p_{0}$$p_{1}$ values such that the two terms are equal. In the new resulting instance, $I^{\prime}_{1}$, ADPATADPAT increases (as we shift probability from lower valued items to a bigger valued item) and $\operatorname{ALG}$ may only decrease.
Next, we have a series of transformations from $I^{\prime}_{i}$ to $I^{\prime}_{i+1}$ for any $i\in[n-1]$: We increase the value $v_{i+1}$ until either (1a) it equals $v_{i}$ or (1b) choosing item 0 and then item $i+1$ yields the same expected value as choosing item 11. That is: $v_{1}=(1-p_{0})(1+\sum_{j=1}^{i}p_{j}v_{i+1})$.
@@
In case (1b), via an additional simple induction, we get that all of the options of $\operatorname{\sigma_{ALG}}$ (of choosing item 11 only, or choosing item 0 and then item $j\in\left\{1,\ldots,i+1\right\}$) yield the same expected value. The process stops when either an instance of n1n-1 items with the same adaptivity gap is obtained (and can therefore use the induction hypothesis) or when we reach $I^{\prime}_{n}$ and get that all options yield the same expected value, as required.
Case (2): Taking item 0 and then item 11 is the (strictly) best option. In this case, we can decrease $p_{1}$ and increase $p_{0}$ by the same amount until this is no longer the case. From intermediate value theorem there is such a $p_{1}$. The adaptivity gap may only increase due to the change: Let $\operatorname{ADAPT}^{\prime}$$\operatorname{ALG}^{\prime}$ be the expected value of $\operatorname{\sigma_{AD}}$ and $\operatorname{\sigma_{ALG}}$ in the new resulting instance. Let $\Delta=(1-p_{0})(p_{1}-p^{\prime}_{1})v_{1}$ be the difference between $\operatorname{ADAPT}$ and $\operatorname{ADAPT}^{\prime}$, and similarly between $\operatorname{ALG}$ and $\operatorname{ALG}^{\prime}$. Then $\frac{\operatorname{ADAPT}^{\prime}}{\operatorname{ALG}^{\prime}}=\frac{\operatorname{ADAPT}-\Delta}{\operatorname{ALG}-\Delta}\geq\frac{\operatorname{ADAPT}}{\operatorname{ALG}}$; the adaptivity gap has only increased. After the change the option of taking only item 11 and the option of taking item 0 and then item 11 both yield the same expected value, which is the value of $\operatorname{ALG}$. We thus may continue as in case (1).
Case (3): If the (strictly) best option of $\operatorname{\sigma_{ALG}}$ is to take item 0 and then some item $i>1$, then $v_{1}<(1-p_{0})(1+\sum_{j=1}^{i}p_{j}v_{i})$. We could decrease $p_{i}$ and increase $p_{1}$ by the same amount until finally either (a) $p_{i}=0$ (and therefore no optimal policy, adaptive or not, will choose item ii) or (b) via intermediate value theorem we get that the option of choosing item 0 and then item 11 yields the same value as the option of choosing item 0 and then item ii.
@@
Case (3b): In the second case, the two options yield the same expected value (while the adaptivity gap may only increase). We continue this way for all $i>1$ so there is no $i>1$ such that taking it after item 0 yields the strictly best option of $\operatorname{\sigma_{ALG}}$. In the resulting instance: either taking only item 11 or taking item 0 and then taking item 11 is the best option. If taking only item 11 is strictly the better option, and we can continue according to case (1) or case (2).